# Surface Reconstruction Using Rotation Systems

### v1

@@ 6. Discussion
Our work is based on the simple observation that a tree connecting all points in a point cloud is a planar graph and, given a rotation system, we can construct a genus zero mesh with a single face that contains all points. The edges of that face are precisely the edges of the tree covered twice, and our results demonstrate that it is possible to triangulate this face by incrementally inserting edges. We can raise the genus by explicitly and judiciously adding handles.
The proposed method outputs a surface reconstruction where all inlier points that belong to the graph that we construct initially are also part of the output mesh. Many other methods reject points in order to construct the best possible output mesh, but in effect this means that reconstruction and noise reduction are subtly coupled. If this is desired it seems that volumetric methods might be more appropriate. In contrast, our method rejects no points and leaves it to the downstream processing to reduce noise. That can be beneficial for instance in applications where several sub-scans are combined, and it is desirable that there is no bias in how many point are kept from each sub-scan.
Compared to previous work, we initially make sure that connected vertices belong to the same face. This allows us to guarantee that gratuitous handles are not added. Of course, the true genus of the reconstructed object is not necessarily known. Hence, it is not generally possible to guarantee correct topology, but since we add handles explicitly, we can impose a a threshold on the handle quality.
@@ Acknowledgements.
Thanks where due.

### v2

@@ 6. Discussion
Our work is based on the simple observation that a tree connecting all points in a point cloud is a genus 0 graph and, given a rotation system, we can construct a mesh with a single face that contains all points. The edges of that face are precisely the edges of the tree covered twice, and our results demonstrate that it is possible to triangulate this face by incrementally inserting edges. We can raise the genus by explicitly and judiciously adding handles.
The proposed method outputs a surface reconstruction where nearly all points that belong to the graph that we construct initially are also part of the output mesh. Many other methods reject a large number of points in order to construct the best possible output mesh, but in effect this means that reconstruction and noise reduction are subtly coupled. If this is desired it seems that volumetric methods might be more appropriate. In contrast, our method only rejects very few points and leaves it to the downstream processing to reduce noise. That can be beneficial for instance in applications where several sub-scans are combined, and it is desirable that there is no bias in how many point are kept from each sub-scan.
Compared to previous work, we initially make sure that connected vertices belong to the same face. This allows us to guarantee that gratuitous handles are not added. Of course, the true genus of the reconstructed object is not necessarily known. Hence, it is not generally possible to guarantee correct topology, but since we add handles explicitly, we can impose a threshold on the handle quality.
@@ Acknowledgements.
We thank all data providers: the Stanford Computer Graphics Laboratory; Qingnan Zhou and Alec Jacobson; Chu et al.; Armeni et al.; Matterport, Inc.; Knapitsch et al.; and Anderson Winkler. We thank the anonymous reviewers for their helpful comments. This work is partially supported by a DTU alliance scholarship, the Danish Council for Independent Research (6111-00552B), the Carlsberg Foundation (CF21-0302), and the German Research Foundation (Gottfried Wilhelm Leibniz programme).